OpenStudy (anonymous):

Wilson is 6 years older than Mary. Wilson's age is 10 years less than five times Mary's age. The equations below model the relationship between Wilson's age (w) and Mary's age (m): w = m + 6 w = 5m − 10 Which is a possible correct method to find Wilson's and Mary's ages?

OpenStudy (anonymous):

@princeharryyy can you help?

OpenStudy (anonymous):

can you help? @dan815

OpenStudy (dan815):

w = m + 6 w = 5m − 10 m + 6 = 5m − 10

OpenStudy (dan815):

m + 6 = 5m − 10 m + 6 -m = 5m - 10 -m 6=4m-10 6+10=4m - 10 + 10 16 = 4m 4=m

OpenStudy (anonymous):

thank you so muchhhh(((: @dan815

OpenStudy (dan815):

m=4 w=m+6 w=4+6=10